A woman invests money in two funds that yield an annual total income of $86. One fund pays an annual return of 7%, and the other pays a return of 8%. How much is invested in the fund yielding 7% if the total annual amount invested at the start of the year is $1200.00
Answer by lwsshak3(11628)

A woman invests money in two funds that yield an annual total income of $86. One fund pays an annual return of 7%, and the other pays a return of 8%. How much is invested in the fund yielding 7% if the total annual amount invested at the start of the year is $1200.00
**
let x=amt invested in 7% fund
1200-x=amt invested in 8% fund
..
.07x+.08(1200-x)=86
.07x+96-.08x=86
.01x=10
x=1000
1200-x=200
amt invested in 7% fund=$1000